OpenStudy (anonymous):

The volume of a pyramid with a rectangular base is V= 1/3 LWH. Find W if V = 500, L = 15, and H = 4.

OpenStudy (anonymous):

Okay, first i would solve for W. Can you do that?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you have to solve this formula for W: 

OpenStudy (anonymous):

\[V=1/3 L W H \]\[\frac{V}{1/3 L\text{ }H}=\frac{1/3 L W H}{1/3 L\text{ }H} \]\[\frac{V}{1/3 L\text{ }H}=\text{ }W \]\[\left(\frac{3 V}{H L}\right)=\text{ }W \]\[\left(\frac{3*500}{4*15}\right)=\text{ }W \]\[(25)=\text{ }W \]